# Closing a catenary loop: the lariat chain, the string shooter, and the heavy elastica

### v2

@@ 5 Some misconceptions
With all the above noted, let us explicitly identify a few misconceptions and some misleading language surrounding this problem that repeatedly appear in the published scientific literature or academic discussions around the IPT: [^4] Regarding additional extensive discussions online on video sharing websites or other social media by professional science popularizers, companies profiting from a redesign and renaming of the original invention, or enthusiastic members of the public, suffice it to say that “someone is wrong on the internet” xkcd .
@@
    The body is not “lifted” by drag forces; it is supported by the force supplied at the launch point. There are no lift (normal drag) forces on a purely tangentially moving body, and the integral of a constant-magnitude tangentially-directed force around a closed loop is zero. However, as has been noted in Gregory49 , Taberlet19 , Daerr19 , there is a moment about the support due to drag, which will balance the moment due to gravity. But while these moments increase if the shape “lifts” to the side, they decrease again if the shape “lifts” farther upwards. While drag can “raise” the shape somewhat, a quite significant effect is achieved by changing the launch angle, as will be shown below in Section 6.1.555We note that visual demonstrations of the necessity of air have been performed with boundary conditions that will not capture such effects clearly, namely either close to horizontal Taberlet19 or vertical Daerr19 . The latter is compared with a lifted-up state that does not appear to be steady; Gregory Gregory49 points out that a vertical boundary condition requires an infinite tension. Axial motion generates both inertial forces, which change the tension, and tangential drag forces, which change the tension and the shape. The former can, for example, stabilize very tall unclosed catenaries MillerYouTube in the absence of air, or make heavy chains hover slowly Aitken1878 , LariatChain . Both this stabilization and the drag-induced bifurcations can be misinterpreted as a sudden lifting action by the air at some threshold speed.
@@ 6 A few solutions and non-solutions
In this section, we present a selection of string and beam solutions to the axially moving loop problem posed above in Section 2. Some of these “solutions” are, like some previously published numerical results, not physically admissible, as they require external forces not present in the problem as posed. We refrain from comment on the stability of the solutions, noting simply that strings under compression ($\sigma<0$) are unstable, but may be stabilized either by adding inertial forces— independently of drag forces— to bring them under tension, or by adding bending stiffness, such that compression is permitted. Inertia alone is irrelevant to most aspects of the problem, serving merely to shift the tension by a constant, whereas drag forces can affect both the shape, and qualitative features of the tension profile.
@@ 6.1 Catenary loops
enforced by a multiplier $\sigma(s,t)$ which we identify as the (bare, unshifted) tension. There is no need to independently consider angular momentum balance. Noting that the curve will be planar, we further define a $\bm{\hat{e}}_{1}$ axis perpendicular to gravity, and a tangential angle $\theta(s,t)$ describing the orientation of the curve’s unit tangent $\bm{\hat{e}}_{t}=d_{s}\bm{x}=\bm{\hat{e}}_{1}\cos\theta+\bm{\hat{e}}_{2}\sin\theta$ and unit normal $\bm{\hat{e}}_{n}=-\bm{\hat{e}}_{1}\sin\theta+\bm{\hat{e}}_{2}\cos\theta$, so that $d_{s}\bm{\hat{e}}_{t}=d_{s}\theta\bm{\hat{e}}_{n}$ and $d_{s}\bm{\hat{e}}_{n}=-d_{s}\theta\bm{\hat{e}}_{t}$. The out of plane unit vector is $\bm{\hat{e}}_{3}$, and both Cartesian and curve frames are right-handed.
@@
where $\bf{R}$} is a singular source of linear momentum, as may be provided only at the support point. In other words, when patching two solutions together, a jump in the quantity $\left(\sigma-v^{2}\right)\bm{\hat{e}}_{t}$ requires an external force. It follows that the shifted tension $\sigma-v^{2}$ must vanish at an unsupported (${\bf{R}}=\bm{0}$) kink. [^6] The exceptional case of a cusp, where $\bm{\hat{e}}_{t}$ jumps by $\pi$, will not be considered.
@@
Figure 2 presents one example each of low, moderate, and high drag catenaries, launched at an angle of $\frac{\pi}{4}$. The low drag string is not a proper solution, as it does not satisfy linear momentum balance without the inclusion of a second force at the vertical point, in addition to the allowed support force at the origin. The inset in (c) reveals the tendency for the tension to diverge at the vertical ($\theta=-\frac{\pi}{2}$). This divergence, and the approach of curvature to zero, are truncated by the presence of a kink, revealed as a gap in angle in the insets in (b) and (d). This gap can only be closed as the length of the string diverges; note that arc length is shown rescaled by this total length. The tension divergence is eliminated at moderate drag, for which subtle but important changes in shape are present, including the curvature curve in (d) reaching the axis at $\theta=-\frac{\pi}{2}$ by orthogonal approach. Further perspective on the derivatives of the $\theta$ curve in (b), which are all zero or infinite at the vertical, can be found in the footnote7. A qualitative change in shape is apparent at high drag, for which the curvature diverges at the vertical, although there is no clear visual signature of the singularity itself.
Figure 3 presents moderate drag catenaries with different launch angles. It can be seen that varying the launch angle “lifts” the string as much as drag does. The solutions are non-unique, as the curves in (d) intersect on the axis at $\theta=-\frac{\pi}{2}$.
@@ 6.2 Regularization by bending stiffness
It is easy enough to physically hang a static loop from a support, or launch a slow-moving one, yet the physics of a catenary forbids such solutions. The most straightforward resolution of this issue comes from the addition of a bending stiffness, which allows for inflectional and vertical points. Such points are clearly present in experiments, including for example in [Daerr19, , figure 1c] or [Taberlet19, , figure 2a], the latter reminiscent of [WatsonWang81, , figure 4d] and [Aitken1878, , plate III figure 3]. The resulting problem is that of the heavy elastica or stiff catenary Wang86 . In the static limit, the physical situation is that of holding an elastic loop at one point, as between two fingers that supply a force and moment. We will also still allow for the presence of a kink.
@@
is potentially redundant, as it provides the same information as the normal projection of the linear momentum balance. However, we use it in our numerical formulation (detailed in A.4), and we will refer to it later in a discussion of global balances (Section 6.3). Defining the quantity appearing in the derivative in (9) as $\bm{f}\equiv\bm{n}-v^{2}\bm{\hat{e}}_{t}$, the shifted tension is now $\bm{f}\cdot\bm{\hat{e}}_{t}=\sigma-v^{2}+E(d_{s}\theta)^{2}$. Jump conditions OReilly17book for these equilibrium configurations are ${\bf{R}}+\llbracket\bm{f}\rrbracket=\bm{0}$ and ${\bf{M}}+\llbracket\bm{m}\rrbracket=\bm{0}$, where the latter condition, allowing for the presence of a singular source of angular momentum $\bf{M}$} at the support, is identical to a static boundary condition for this beam model without rotary inertia. The only jump is at the support point, as there is no longer a need for patching at the vertical. It is apparent from from (10) that the $d_{s}^{3}\theta$ term, the second derivative of curvature, is what will balance the product of shifted tension and cuvature at the vertical, where $\cos\theta=0$.
Figure 4: Stiff catenary loops with bending stiffness $E=0.00254$ and zero, low, and moderate drag ($D=0,0.90,1.10$) and a kink-free vertical orientation at the support (launch angle $\theta(0)=\frac{\pi}{2}$, return angle $\theta(1)=-3\frac{\pi}{2}$), alongside a classical static non-stiff ($D=0,E=0$) catenary with an angle cutoff of $\pm 0.05$. (a) Shapes. (b) Arc length $s$, (c) shifted tension $\sigma-v^{2}+E\left(d_{s}\theta\right)^{2}$, and (d) curvature $d_{s}\theta$, as functions of angle $\theta$. The curvature of the catenary is approximately 79.933-79.933 at $\pi$ and 0, and approximately 0.200-0.200 at the patching point.
@@
Figure 4 presents stiff catenaries (beams) with zero, low, and moderate drag, and a kink-free vertical orientation at the support, alongside a static non-stiff catenary (string) with small angle cutoff for qualitative comparison. The catenary fails to satisfy momentum balance by a large amount. Its angle gap, diverging tension, and nonzero curvature at the patching point near the vertical $\theta=-\frac{\pi}{2}$ are clearly visible across (b), (c), and (d). These features are absent in the presence of bending stiffness. No obvious signature of the $D=1$ bifurcation remains. Both the shifted tension and the curvature are nonzero at the vertical.
Figure 5 presents stiff catenaries with moderate drag, launch angle $\frac{\pi}{4}$ and different return angles or stiffnesses, alongside a non-stiff catenary with the same drag, launch angle, and return angle as two of the beams. Allowing a kink in an axially moving elastic beam can instead be interpreted as material being fed in and out at the launch point, rather than a driven closed loop. Note that both inflectional and non-inflectional elastica are represented, depending on boundary conditions.
It is clear that even a small amount of bending stiffness leads to drastic changes in shape, creating the eggplant or slumping balloon shapes of Figures 4-5.
@@ 6.3 Global balances
the latter also following directly from the definition of $\bm{m}$. These two quantities can be identified as $\bf{R}$} and $\bf{M}$} at the support point, where a jump $\llbracket q\rrbracket=q(0)-q(1)$. The force is purely vertical, balancing gravity; the drag does not contribute. In the string limit ($E=0$), it is not possible to apply a moment, and as pointed out in several previous works Gregory49 , Taberlet19 , Daerr19 , the integral moments about the support due to drag and gravity balance each other; we can also infer that the moment about the center of mass due to drag balances that due to the support force.
@@ 7 Open questions
The possibility of other forms of solutions, such as undulatory shapes obtained by patching in the moderate drag regime, has not been considered.
The vertical singularity, at which derivatives of angle either vanish or blow up, would be interesting to explore further, particularly in the moderate drag range where all the transitions of the derivatives occur, as discussed in a previous footnote.7 The sequence of shape bifurcations is likely to be visually subtle, but may lead to numerical issues. We are not immediately aware of another simple physical problem generating a toy mathematical problem with this type of singular behavior. Whether any transitions remain or arise upon the addition of bending is unknown.
@@ 8 Summary and Conclusions
We have corrected several errors in recent works on axially moving string loops in the presence of gravity. These entail misunderstandings of the role of drag forces, bifurcations, the nature of the vertical singularity, and admissibility of certain solutions. Several example solutions were constructed, either by implementing previously derived analytical results Gregory49 , ChakrabartiHanna16 , or numerically adding bending stiffness. We have also attempted to place the problem in context alongside several other fundamental problems in flexible structure mechanics.

### v3

@@ 5 Some misconceptions
After all the above, let us briefly note a few recurring misconceptions and some misleading language surrounding this problem that appear in the published scientific literature or academic discussions around the Tournament problem: [^5] Regarding additional extensive discussions online on video sharing websites or other social media by professional science popularizers, companies profiting from a redesign and renaming of the original invention, or enthusiastic members of the public, suffice it to say that “someone is wrong on the internet” xkcd .
@@
    The body is not “lifted” by drag forces; it is supported by the force supplied at the launch point. As noted in Kurkin64 , Taberlet19 , Daerr19 , there are no lift (normal drag) forces on a purely tangentially moving body, and the integral of a constant-magnitude tangentially-directed force around a closed loop is zero. However, as noted in Gregory49 , Kurkin64 , Miroshnik01 , Taberlet19 , Daerr19 , there is a moment about the support due to drag, which will balance the moment due to gravity. But while these moments increase if the shape “lifts” to the side, they decrease again if the shape “lifts” farther upwards. While drag can “raise” the shape somewhat, a quite significant effect is achieved by changing the launch angle, as was shown for high drag in Miroshnik01 (see also Miroshnik72inrussian for a case with additional normal flow), and will be shown below for low drag in Section 6.1.666We note that recently published visual demonstrations of the necessity of air have been performed with boundary conditions either close to horizontal Taberlet19 or vertical Daerr19 . The latter is compared with a lifted-up state that does not appear to be steady; Gregory Gregory49 points out that a vertical boundary condition requires an infinite tension. Axial motion generates both inertial forces, which change the tension, and tangential drag forces, which change the tension and the shape. The former can, for example, stabilize very tall unclosed catenaries MillerYouTube in the absence of air, or make heavy chains hover slowly Aitken1878 , LariatChain . Both this stabilization and the drag-induced bifurcations could resemble a sudden lifting action by the air at some threshold speed.
@@ 6 A few solutions and non-solutions
In this section, we present a selection of string and beam solutions to the axially moving loop problem posed above in Section 2. Some of these “solutions” are, like some previously published numerical results, not physically admissible, as they require external forces not present in the problem as posed. We do not examine the stability of solutions, and do not anticipate that this will display any interesting complexity, noting simply that strings under compression ($\sigma<0$) are unstable, but may be stabilized either by adding inertial forces— independently of drag forces— to bring them under tension, or by adding bending stiffness, such that compression is permitted. Inertia alone is irrelevant to most aspects of the problem, serving merely to shift the tension by a constant, whereas drag forces can affect both the shape, and qualitative features of the tension profile.
@@ 6.1 Catenary loops
enforced by a multiplier $\sigma(s,t)$ which we identify as the (bare, unshifted) tension. There is no need to independently consider angular momentum balance. Noting that the curve will be coplanar with gravity, we further define a $\bm{\hat{e}}_{1}$ axis perpendicular to gravity, and a tangential angle $\theta(s,t)$ describing the orientation of the curve’s unit tangent $\bm{\hat{e}}_{t}=d_{s}\bm{x}=\bm{\hat{e}}_{1}\cos\theta+\bm{\hat{e}}_{2}\sin\theta$ and unit normal $\bm{\hat{e}}_{n}=-\bm{\hat{e}}_{1}\sin\theta+\bm{\hat{e}}_{2}\cos\theta$, so that $d_{s}\bm{\hat{e}}_{t}=d_{s}\theta\bm{\hat{e}}_{n}$ and $d_{s}\bm{\hat{e}}_{n}=-d_{s}\theta\bm{\hat{e}}_{t}$. The out of plane unit vector is $\bm{\hat{e}}_{3}$, and both Cartesian and curve frames are right-handed.
@@
where $\bf{R}$} is a singular source of linear momentum, as may be provided only at the support point. In other words, when patching two solutions together, a jump in the quantity $\left(\sigma-v^{2}\right)\bm{\hat{e}}_{t}$ requires an external force. It follows that the shifted tension $\sigma-v^{2}$ must vanish at an unsupported (${\bf{R}}=\bm{0}$) kink. [^7] The exceptional case of a cusp, where the orientation of $\bm{\hat{e}}_{t}$ jumps by $\pi$, will not be considered.
@@
Figure 2 presents one example each of low, moderate, and high drag catenaries, launched at an angle of $\frac{\pi}{4}$. The low drag string is not a proper solution, as it does not satisfy linear momentum balance without the inclusion of a second force at the vertical point, in addition to the allowed support force at the origin. The inset in (c) reveals the tendency for the tension to diverge at the vertical ($\theta=-\frac{\pi}{2}$). This divergence, and the approach of curvature to zero, are truncated by the presence of a kink, revealed as a gap in angle in the insets in (b) and (d). This gap can only be closed as the length of the string diverges; note that arc length is shown rescaled by this total length. The tension divergence is eliminated at moderate drag, for which subtle but important changes in shape are present, including the curvature curve in (d) reaching the axis at $\theta=-\frac{\pi}{2}$ by orthogonal approach. Further perspective on the derivatives of the $\theta$ curve in (b), which are all zero or infinite at the vertical, can be found in the footnote8. A qualitative change in shape is apparent at high drag, for which the curvature diverges at the vertical, although there is no clear visual signature of the singularity itself.
@@
Figure 3 presents moderate drag catenaries with different launch angles. It can be seen that varying the launch angle “lifts” the string as much as drag does. The solutions are non-unique, as the curves in (d) intersect on the axis at $\theta=-\frac{\pi}{2}$. However, this seems unlikely to correspond to any possibility of coexisting states in the string shooter problem as typically posed. One might fix the drag $D$ by setting the speed, and impose some launch angle. The other features of the response appear to be unique. For example, the return angle would be determined by Gregory’s Gregory49 transcendental relation (25), reproduced in our A.3.
@@ 6.2 Regularization by bending stiffness
It is easy enough to physically hang a static loop from a support, or launch a slow-moving one, yet the physics of a catenary forbids such solutions. The most straightforward resolution of this issue comes from the addition of a bending stiffness, which allows for nonvertical inflectional, and vertical, points. Such points are clearly present in experiments, including for example in [SvetlitskiiGabryuk66, , figure 3], [Miroshnik01, , figures 10 and 11], [Daerr19, , figure 1c], and [Taberlet19, , figure 2a], the latter reminiscent of [WatsonWang81, , figure 4d] and [Aitken1878, , plate III figure 3]. The resulting problem is that of the heavy elastica or stiff catenary Wang86 . In the static limit, the physical situation is that of holding an elastic loop at one point, as between two fingers that supply a force and moment. We will also still allow for the presence of a kink.
@@
is potentially redundant, as it provides the same information as the normal projection of the linear momentum balance. However, we use it in our numerical formulation (detailed in A.4), and we will refer to it later in a discussion of global balances (Section 6.3). Defining the quantity appearing in the derivative in (9) as $\bm{f}\equiv\bm{n}-v^{2}\bm{\hat{e}}_{t}$, the shifted tension is now $\bm{f}\cdot\bm{\hat{e}}_{t}=\sigma-v^{2}+E(d_{s}\theta)^{2}$. Jump conditions OReilly17book for these equilibrium configurations are ${\bf{R}}+\llbracket\bm{f}\rrbracket=\bm{0}$ and ${\bf{M}}+\llbracket\bm{m}\rrbracket=\bm{0}$, where the latter condition, allowing for the presence of a singular source of angular momentum $\bf{M}$} at the support, is identical to a static boundary condition for this beam model without rotary inertia. The only jump is at the support point, as there is no longer a need for patching at the vertical. It is apparent from (10) that the $d_{s}^{3}\theta$ term, the second derivative of curvature, is what will balance the product of shifted tension and curvature at the vertical, where $\cos\theta=0$.
Figure 4: Stiff catenary loops with bending stiffness $E=0.00254$ and zero, low, and moderate drag ($D=0,0.90,1.10$) and a kink-free vertical orientation at the support (launch angle $\theta(0)=\frac{\pi}{2}$, return angle $\theta(1)=-\frac{3\pi}{2}$), alongside a classical static non-stiff ($D=0,E=0$) catenary with an angle cutoff of $\pm 0.05$. (a) Shapes. (b) Arc length $s$, (c) shifted tension $\sigma-v^{2}+E\left(d_{s}\theta\right)^{2}$, and (d) curvature $d_{s}\theta$, as functions of angle $\theta$. The curvature of the catenary is approximately 79.933-79.933 at $\pi$ and 0, and approximately 0.200-0.200 at the patching point.
Figure 4 presents stiff catenaries (beams) with zero, low, and moderate drag, and a kink-free vertical orientation at the support, alongside a static non-stiff catenary (string) with a small angle cutoff for qualitative comparison. The catenary fails to satisfy momentum balance by a large amount. Its angle gap, diverging tension, and nonzero curvature at the patching point near the vertical $\theta=-\frac{\pi}{2}$ are clearly visible across (b), (c), and (d). These features are absent in the presence of bending stiffness. No obvious signature of the $D=1$ bifurcation remains. Both the shifted tension and the curvature are nonzero at the vertical.
@@
Figure 5 presents stiff catenaries with moderate drag, launch angle $\frac{\pi}{4}$, and different return angles or stiffnesses, alongside a non-stiff catenary with the same drag, launch angle, and return angle as two of the beams. Allowing a kink in an axially moving elastic beam can instead be interpreted as material being fed in and out at the launch point, rather than a driven closed loop. Note that both inflectional and non-inflectional elastica are represented, depending on boundary conditions.
It is clear that even a small amount of bending stiffness leads to drastic changes in shape, creating the eggplant or slumping balloon shapes of Figures 4-5. Figure 5(a) can be qualitatively compared with the moderate drag experimental shapes from [SvetlitskiiGabryuk66, , figure 3] and [Miroshnik01, , figure 11], although these works do not consider bending.
@@ 6.3 Global balances
the latter also following directly from the definition of $\bm{m}$. These two quantities can be identified as $\bf{R}$} and $\bf{M}$} at the support point, where a jump $\llbracket q\rrbracket=q(0)-q(1)$. The force is purely vertical, balancing gravity; the drag does not contribute. In the string limit ($E=0$), it is not possible to apply a moment, and as noted in Gregory49 , Taberlet19 , Daerr19 , the integral moments about the support due to drag and gravity balance each other; we can also infer that the moment about the center of mass due to drag balances that due to the support force.
@@ 7 Open questions
The vertical singularity, at which derivatives of angle either vanish or blow up, would be interesting to explore further, particularly in the moderate drag range where all the transitions of the derivatives occur, as discussed in a previous footnote.8 The sequence of shape bifurcations is likely to be visually subtle, but may lead to numerical issues. We are not immediately aware of another simple physical problem generating a toy mathematical problem with this type of singular behavior. Whether any transitions remain or arise upon the addition of bending is unknown.
@@ 8 Summary and Conclusions
We have corrected several errors in recent works on axially moving string loops coplanar with gravity. These entail misunderstandings of the role of drag forces, bifurcations, the nature of the vertical singularity, and admissibility of certain solutions. Several example solutions were constructed, either by implementing previously derived analytical results, or numerically adding bending stiffness. We have also attempted to place the problem in context alongside several other fundamental problems in flexible structure mechanics.